### v1

# Beyond Wide-Angle Images: Unsupervised Video Portrait Correction via Spatiotemporal Diffusion Adaptation

Abstract
Wide-angle cameras, despite their popularity for content creation, suffer from distortion-induced facial stretching—especially at the edge of the lens—which degrades visual appeal. To address this issue, we propose an image portrait correction framework using diffusion models named ImagePD. It integrates the long-range awareness of transformer and multi-step denoising of diffusion models into a unified framework, achieving global structural robustness and local detail refinement. Besides, considering the high cost of obtaining video labels, we then repurpose ImagePD for unlabeled wide-angle videos (termed VideoPD), by spatiotemporal diffusion adaption with spatial consistency and temporal smoothness constraints. For the former, we encourage the denoised image to approximate pseudo labels following the wide-angle distortion distribution pattern, while for the latter, we derive rectification trajectories with backward optical flows and smooth them. Compared with ImagePD, VideoPD maintains high-quality facial corrections in space and mitigates the potential temporal shakes sequentially. Finally, to establish an evaluation benchmark and train the framework, we establish a video portrait dataset with a large diversity in people number, lighting conditions, and background. Experiments demonstrate that the proposed methods outperform existing solutions quantitatively and qualitatively, contributing to high-fidelity wide-angle videos with stable and natural portraits. The codes and dataset will be available.
Figure 1: Our wide-angle video portrait correction results. (a) The first row shows a wide-angle phone video with edge distortion, corrected in the second row using our method while preserving spatiotemporal consistency. (b) Video correction introduces temporal shake (orange line in the figure), while our method smooths the trajectory (blue line).
1 Introduction
With the development of the self-media and videography industries, wide-angle lenses have become increasingly favored for the capability of capturing expansive scenes. However, such lenses inevitably introduce geometric distortions, particularly pronounced at the lens boundaries, resulting in background straight line curvatures and facial feature deformations in still images and video recordings.
Traditional wide-angle portrait correction methods typically require precise camera parameters (e.g., focal length [29]) as a prerequisite, followed by stereographic projection, face and line detection, and energy optimization to achieve geometric correction [29]. Such a pipeline is complex in process and inapplicable when the relevant camera parameters are unknown. The same goes for video portrait correction work [17]. In contrast, learning-based solutions eliminate this issue [33] by directly learning the spatial mapping from wide-angle images to rectification flows in a supervised manner. However, these methods often exhibit noticeable temporal shakes when applied to videos. Moreover, there is currently a lack of datasets for video portrait correction due to the significantly higher cost of labeling video data.
To this end, we take a pioneering step in achieving video portrait correction with diffusion models. First of all, we present a global-to-local portrait correction diffusion model, named ImagePD, to generate high-quality rectification flows, which integrates the advantages of both Transformer and diffusion models. The Transformer model establishes long-range geometric dependencies by capturing spatial global features across entire images, providing structural guidance for the subsequent diffusion process. The diffusion model refines high-fidelity flow patterns through iterative denoising guided by Transformer-derived structural information. This co-design enables structure-to-detail pixel-wise correction, where structural priors from the Transformer and detail-generation capacity from the diffusion model are synergistically combined.
More importantly, we propose an unsupervised spatiotemporal diffusion adaption approach to enable ImagePD the capability to rectify video sequences stably. The repurposed model, which we term VideoPD, addresses the issue of temporal shakes without requiring video correction labels. Concretely, we first use the pre-trained ImagePD to generate rectification flows for each video frame as pseudo-labels. When ImagePD is adapted to unseen wide-angle videos, these pseudo-labels can guide the denoising process, ensuring spatial consistency in facial corrections with the wide-angle distortion distribution prior. To enforce temporal smoothness, we derive rectification trajectories across sequential frames using backward optical flows and design a temporal smoothness constraint to mitigate sequential jitters. By unsupervised spatiotemporal adaptation with diffusion models, VideoPD achieves stable wide-angle video correction that not only preserves high-fidelity facial details but also frees from labor-intensive video annotations.
To establish an evaluation benchmark and train the proposed video portrait correction model, we construct a wide-angle video dataset with a wide diversity in scene, camera, and number of people. Finally, we conduct extensive experiments about portrait correction in both image and video, demonstrating our superiority over other solutions. The contributions center around:
  • We propose the first portrait correction diffusion model, named ImagePD, which integrates long-range dependences of transformer and iterative refinements of diffusion models for global-to-local rectification.
  • We design an unsupervised spatiotemporal diffusion adaption framework to transfer ImagePD to VideoPD without labeled videos. It combines spatio-temporal optical flows to track correction trajectories, establishing spatial consistency and temporal smoothness constraints.
  • We present a video portrait dataset with wide-angle cameras and conduct extensive experiments to validate our superior performance over existing solutions.
2 Related Work
2.1 Portrait Correction
Portrait correction in wide-angle images is a significant research area. Traditional methods, such as Shih et al.[29]’s content-aware warping, adjust distortions by analyzing image content and camera parameters. However, these methods bbnnv nbrely on camera parameters or complex inputs and are difficult to extend to video processing. Stereographic projection methods[29] can preserve local conformality but struggle with temporal consistency in videos.
Deep learning has introduced calibration-free solutions. Tan et al.[33] proposed a two-stage neural network leveraging wide-angle images and correction flow datasets. In contrast, Zhu et al. [47] developed a semi-supervised Transformer to reduce annotation costs. Nie et al.’s method CoupledTPS[23], iteratively couples limited TPS models with a warping flow to precisely correct facial features.
Portrait correction in wide-angle videos has garnered less research focus than image-based studies. Current deep learning methods, designed solely for static image processing, do not accommodate the temporal dynamics inherent in video data. This stems from the high cost of manually annotating wide-angle videos, which demands frame-by-frame effort and proves impractical for continuous video tasks, resulting in a scarcity of comprehensive training datasets.
Lai et al.[17]. introduced temporal consistency energy minimization based on traditional spherical projection and energy minimization to optimize videos. However, this method lacks a connection between the global information of wide-angle videos and the detailed information needed for facial correction. It also lacks adaptability and stability when dealing with complex environments and conditions.
To address this issue, we propose learning video correction principles from image networks.
2.2 Video stabilization
Traditional video stabilization techniques primarily stabilize the video by smoothing the feature trajectories between multiple frames or adjacent frames [35, 18, 28]. Subsequently, grid-based methods [35, 20, 22] and optical flow (Liu et al.[21]; Geo et al.[15]) have also been applied to represent motion, becoming key tools for stabilizing videos. With the rapid development of deep learning, deep learning methods have been applied to video stabilization technology [45, 43, 46, 25, 39], where the original video frames are input, and stabilized video frames are directly output. For example, Wang et al. [34] proposed an end-to-end learning framework, and Xu et al. [37] used adversarial networks to generate target images.DUT [38], which learns video stabilization by simply watching unstable videos. Additional methods, including unsupervised online video stitching (Nie et al.[24]), and meta-learning-based approaches to improve full-frame stabilization (Ali et al.[1]), have been explored to enhance stability. These approaches, ranging from optical flow to deep learning models, provide valuable solutions for addressing video temporal shake in video portrait correction.
2.3 Diffusion Models
Diffusion models (DMs)[40] have recently achieved remarkable success in generative tasks, extending from image generation to video-related applications. In video generation, early advancements include Video Diffusion Models (VDM)[13], which adapted the 2D U-Net architecture from image DMs to a 3D U-Net for generating coherent video frames. To enhance temporal resolution, Ho et al.[12] introduced a cascade of video super-resolution models, improving the quality of generated sequences. Additionally, methods like Make-A-Video [30] achieved text-to-video generation by leveraging pre-trained text-to-image DMs, avoiding the need for paired text-video data.
More recent work focuses on video editing and enhancing temporal consistency. For example, Tune-A-Video [36] fine-tunes models on single input videos to produce coherent video outputs, while FateZero [26] ensures temporal consistency using attention-based blending techniques. Similarly, Rerender-A-Video [41] and VideoControlNet[14] introduce pixel-aware latent fusion and motion-aware adjustments to maintain visual consistency during edits.
Figure 2: An overview of the proposed framework for portrait distortion correction, integrating Transformer networks and diffusion models. The framework consists of two stages: ImagePD for generating high-quality rectification flow and VideoPD for applying this correction to videos, ensuring spatiotemporal consistency and smooth stabilization.
3 Methodology
3.1 Overview
The overview of our method is shown in Fig. 2, where our framework operates in two cascaded stages: correction and stabilization.In the first stage, ImagePD tackles the distorted image $I_{s}$, producing precise optical flows using a Transformer-guided diffusion approach. The refined optical flow is then applied to generate a geometrically aligned image $\hat{I_{s}}$.
In the second stage, temporal pixel trajectories $T$ are extracted from shaky video sequences, represented as the unstable video frames $V_{s}$. Using the corrected frames $\hat{I}_{s}$ as the reference, the final corrected and stabilized output $V_{\text{final}}$ is computed as:
$V_{\text{final}}=C(V_{s}\mid\hat{I_{s}},T),$ (1)
where $C$ denotes the spatiotemporal fusion module.
3.2 ImagePD
As illustrated in Fig. 2(a,b), our ImagePD is a framework for high-quality portrait correction by predicting precise optical flow fields to rectify facial distortions.
3.2.1 Model Architecture
In the first stage of ImagePD, we focus on capturing the global geometric structure of the distorted input image $I_{s}$. The Transformer network [47], denoted as MS-Unet in Fig. 2(a), with its advantage in capturing long-range information, extracts a set of intermediate layer features $h$ that effectively reflect the global geometric structure of the distorted input image. These features $h$ serve as a structural prior, guiding the diffusion model in the next stage to accurately align the facial geometry while preserving spatial coherence.
Building on the global structure from the first stage, this stage, as shown in Fig. 2(b), fines local details using a diffusion-based approach.
In order to ensure that the correction process primarily focuses on the area of the face, we use InsightFace [27, 2, 3, 6, 5, 7, 8, 10, 11, 9] as a pre-trained model to process the original facial image $I_{\text{s}}$, allowing accurate face detection and keypoint localization. Based on these detection results, a binary mask M is generated, which clearly delineates the facial region. This mask provides spatial guidance for subsequent processing and improves the quality of the correction. Leveraging the detail-handling capability of diffusion models, we form the conditioning input for ImagePD by concatenating $M$, $I_{s}$, and $h$:
$x_{in}=cat[M,I_{s},h]$ (2)
Our approach leverages the iterative refinement capabilities of Denoising Diffusion Implicit Models (DDIM)[31], enabling iterative data generation through noise-adding and denoising processes. In DDIM, an original sample $x_{0}$ is gradually corrupted with Gaussian noise over $T$ timesteps, eventually approximating a standard Gaussian distribution $x_{T}\sim\mathcal{N}(0,I)$. The reverse process starts from $x_{T}$ and iteratively removes noise to recover a clean sample, modeled as a learned denoising transition.
A time-conditioned U-Net, denoted as $\epsilon_{\theta}$, is trained to predict the noise $\epsilon$ added at each step. During training, noise is added to $x_{0}$ to generate intermediate states $x_{t}$, and the model optimizes the following mean squared error (MSE) loss:
$L=\mathbb{E}_{t\sim U(1,T),\,x_{0}\sim q(x_{0}),\,\epsilon\sim\mathcal{N}(0,I)}\Big{[}\|\epsilon-\epsilon_{\theta}(x_{t},t)\|^{2}\Big{]}$ (3)
Adapting this concept, ImagePD redefines the diffusion process to operate directly on optical flow fields.
The task of the diffusion network at this stage is not to directly generate the corrected facial image, but to predict an optical flow field $F^{corr}$. Compared to directly generating an image, the optical flow field offers sub-pixel precision, enabling more fine-grained guidance and correction of minor deformations in the image.
By iteratively refining the predicted optical flow using the diffusion model, our method ultimately produces a highly accurate optical flow prediction. This prediction is then used to warp distortion correction on the facial image, producing the corrected output:
$\hat{I_{s}}=\mathcal{W}(F^{corr},I_{s})$ (4)
The warping operation $\mathcal{W}$ realigns the pixels in $I_{s}$ according to the predicted flow, resulting in a natural facial appearance.
3.2.2 Optimization Objective
The model is optimized using a comprehensive loss function:
$\mathcal{L}_{image}=\lambda_{1}\mathcal{L}_{mask}+\lambda_{2}\mathcal{L}_{photo}+\lambda_{3}\mathcal{L}_{flow}$ (5)
where $\mathcal{L}_{flow}$ ensures optical flow accuracy, $\mathcal{L}_{photo}$ enforces image quality, and $\mathcal{L}_{mask}$ enhances facial edge details, with $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ as weighting coefficients.
Training objective: $\mathcal{L}_{image}$ comprises three components, each addressing a specific aspect of the correction process.
Optical Flow Loss $\mathcal{L}_{flow}$: Measures the difference between the model-predicted optical flow and the ground truth optical flow.
$\mathcal{L}_{flow}=\frac{1}{HW}\sum_{h=1}^{H}\sum_{w=1}^{W}W\left(F^{Corr}(h,w)-F_{\text{gt}}(h,w)\right)^{2}$ (6)
Image Loss $\mathcal{L}_{photo}$: Evaluates the difference between the corrected image after optical flow correction and the ground truth image.
$\mathcal{L}_{photo}=\frac{1}{HW}\sum_{h=1}^{H}\sum_{w=1}^{W}W\left(\hat{I}_{s}(h,w)-I_{\text{gt}}(h,w)\right)^{2}$ (7)
Mask-based Sobel Loss $\mathcal{L}_{mask}$: Uses the Sobel operator to extract edge information and calculates the edge loss within the facial region mask.
$\mathcal{L}_{mask}$ $=\left[\left|G_{x}(F^{Corr})-G_{x}(F_{\text{gt}})\right|\right.$ (8)
$\quad+\left|G_{y}(F^{Corr})-G_{y}(F_{\text{gt}})\right|\big{]}\cdot M$
3.3 VideoPD
Videos are dynamic sequences composed of continuous multi-frame images =$\mathcal{I}=\left\{\mathcal{I}_{1},\mathcal{I}_{2},\dots,\mathcal{I}_{n}\right\}$. To achieve unsupervised wide-angle video correction, we use pre-trained ImagePD to generate corrected optical flow for each frame as pseudo-labels. However, directly applying these pseudo-labels to the video frame sequence for correction results in severe temporal shakes sequentially, significantly decreasing the video’s quality and visual experience. To address this issue, we propose further fine-tuning of ImagePD to enhance its adaptability in distortion video scenarios by introducing VideoPD.
3.3.1 Correction Trajectory
Our method aims to modulate and control the motion between corresponding frames, so we first use the RAFT [44] network to estimate the inter-frame optical flow from the original wide-angle video. For each pair of consecutive frames $I_{t}$ and $I_{t+1}$, RAFT computes the backward optical flow $f_{t\to t+1}$to avoid generating invalid regions. Next, each frame is processed by the ImagePD module, which leverages a diffusion model to predict the optical flow of correction $F^{corr}_{t}$, effectively serving as a pseudo-label. This flow contains detailed adjustment information for the facial regions, addressing distortions introduced by the wide-angle lens.
To address the temporal shake caused by independent corrections between frames, we construct a series of two-dimensional meshes for each frame on the original frames. The pseudo-label optical flow $F^{corr}_{t}$ is then used to adjust the positions of the control points in the mesh and we can obtain an updated mesh $G$.
To ensure smooth corrections across time, the corrected mesh $G$ is used for reprojection onto the optical flow $f_{t\to t+1}$ from RAFT and the pseudo-label flow $F^{corr}_{t}$. This fusion yields a refined temporal optical flow $F^{final}_{t}$, which effectively mitigates warping shakes and ensures consistency between frames. A visual representation of this relationship Eq. 9 is provided in Fig. 3.
$r(t+1)=f_{t\to t+1}\Bigl{(}G+F_{t}^{\mathrm{Corr}}\Bigr{)}+F_{t}^{\mathrm{Corr}}-F_{t+1}^{\mathrm{Corr}}$ (9)
It can be further concatenated in temporal order from the initial moment to obtain the position at each moment as follows:
$R(t)=r(1)+r(2)+\cdots+r(t)$ (10)
where $r(1)$ is defined as a zero matrix. The rectified motion trajectory can be obtained by sequentially concatenating $R(1),R(2),\ldots,R(t),\ldots,R(N)$ over time. We achieve smooth trajectory stabilization and resolve video shaking issues by controlling the incremental stability of this pixel-level dense trajectory.
Figure 3: Illustration of the computation of $r(t+1)$ in Eq. 9, depicting the integration of RAFT optical flow $f_{t\to t+1}$, corrected mesh $G$, and correction flows $F_{t}^{\mathrm{Corr}}$ and $F_{t+1}^{\mathrm{Corr}}$ for temporal stabilization.
3.3.2 Spatiotemporal consistency
To achieve unsupervised video correction, we design our objective function $\mathcal{L}_{video}$ concerning two aspects: spatial consistency and temporal smoothness.
$\mathcal{L}_{video}=\mathcal{L}_{spatial}+\lambda\mathcal{L}_{temporal}$ (11)
For the temporal smoothness, we use the previously derived video warping motion trajectories to achieve this functionality, constraining the video trajectories obtained from the domain-adapted ImagePD model to be as smooth as possible. For three consecutive frames, we encourage the trajectory position of the middle frame to lie between the trajectory positions of the preceding and following frames:
$\mathcal{L}_{temporal}=\|R(t+1)+R(t-1)-2R(t)\|$ (12)
For spatial consistency, we leverage the pre-trained ImagePD model to ensure the quality of facial distortion correction within each frame. This is achieved by incorporating pseudo-labels generated by ImagePD, supervising the optical flow loss $\mathcal{L}_{flow}$ (Eq. 6) and the mask-based Sobel loss $\mathcal{L}_{mask}$(Eq. 8). These components maintain the structural integrity and edge details of the facial regions across individual frames, adapting the spatial correction capabilities from the image domain to the video domain.
3.4 Data preparation
Due to the lack of publicly available wide-angle video portrait datasets, we introduce a new dataset containing 136 clips, each with a resolution of 1080p at 30fps, ranging from 5 seconds to 90 seconds in duration(totaling 43,903 frames). The clips were captured using the iPhone 13 Pro, Gopro Hero 7, Samsung Galaxy S24, and Oneplus ace 2v, covering a wide variety of scenarios such as campuses, parks, shopping malls, classrooms, and other everyday settings. These clips encompass diverse scenarios, including content captured by self-media creators and film clips extracted from cinematic works. The number of people appearing in each scene varies from 1 to 6, reflecting diverse human dynamics. To enhance the dataset’s diversity, we’ve also included footage shot in special conditions like nighttime environments. To authentically replicate real-world recording conditions, we deliberately introduced artificial shaking to simulate the instability inherent in handheld video capture. This design choice poses significant challenges for stabilization algorithms. Fig. 4 presents some examples of our wide-angle video portrait dataset.
4 Experiments
4.1 Dataset and Implement Details
Our methodology builds upon two core datasets: the established wide-angle image dataset [33] and our newly captured video dataset.
Image Data:
In our experiments, we use the dataset from Tan et al.[33], which consists of 5133 training images and 129 test images captured with five wide-angle smartphones.
Figure 4: Our wide-angle video face dataset encompasses a broad range of scenes and showcases significant diversity in individuals. We present some examples from our dataset.
Video Data:
For both training and testing phases, we utilized the dataset described in Sec. 3.4. The training set comprised 16 video sequences (7,134 frames) that comprehensively represent diverse capture devices, environmental conditions, and scene types, ensuring the model training across varied scenarios.
Training Process of ImagePD:
Our framework employs a two-stage training protocol. In the first stage, we utilize the Adam optimizer[16] with an initial learning rate of $1\times 10^{-4}$, training the model for 200 epochs. The subsequent texture refinement stage leverages high-dimensional features extracted from the initial phase. This feature, combined with image $I_{s}$ and mask $M$, acts as the conditional guidance for the diffusion model to enhance fine-grained texture details. This stage implements a learning rate of $2\times 10^{-4}$, with optimization conducted over 500,000 iterations.
Training Process of VideoPD:
For VideoPD, we fine-tune the ImagePD model to maintain correction efficacy while smoothing temporal jitter. Training employs the Adam optimizer[16] with a learning rate of $2\times 10^{-4}$ over 300,000 iterations. The entire training process is conducted on a single RTX 4090.
4.2 Evaluation Metrics
Following Tan et al.’s work  [33], we evaluate geometric correction performance using two established metrics: LineAcc quantifies how well straight edges retain geometric integrity by computing deviations between local slopes and the ideal global orientation. ShapeAcc assesses shape preservation via cosine similarity between reference and corrected landmarks, where higher values indicate better alignment. We also employ Jin et al.’s Stability Score (Avg, Trans, Rot)[4], which uses homography and FFT, to measure VideoPD’s shake reduction, with higher scores denoting smoother videos.
4.3 Comparisons.
We evaluate our approach against existing methods across image and video modalities, providing comprehensive quantitative and qualitative results.
Method Reference LineAcc ShapeAcc
Shih et al.[29] TOG2019 66.143 97.253
Tan et al.[33] CVPR2021 66.784 97.490
Zhu et al.[47] CVPR2022 66.825 97.491
MOWA[19] arXiv 2024 - 97.475
DualPriors[42] ECCV2024 67.304 99.012
CoupledTPS[23] TPAMI2024 66.808 97.500
ImagePD (Ours) - 66.898 97.508
Table 1: Quantitative comparison of the proposed ImgaePD and different portrait correction methods.Higher LineAcc and ShapeAcc indicate better straight-line and face correction performance, respectively.
Figure 5: Comparison of ImagePD with existing methods on wide-angle distorted images. ImagePD achieves superior restoration of natural facial features and reduced edge distortions compared to state-of-the-art alternatives.
Image Comparisons. As shown in Table  1, our ImagePD achieves the second-best performance across LineAcc and ShapeAcc metrics, surpassed only by DualPriors[42]. This shows that our image correction model can retain the geometric integrity of straight edges and preserve facial shape well. Furthermore, in the qualitative comparison presented in Fig. 5, we can further observe that ImagePD outperforms DualPriors[42] in the aforementioned aspects. This undoubtedly demonstrates that, despite DualPriors[42] achieving state-of-the-art performance on certain metrics, their approach exhibits noticeable abrupt artifacts at stitching boundaries and inconsistent head-to-body proportions in the images. This may be attributed to their method relying on a parallel network that separately rectifies the background and portrait before stitching them into a single image, with blank areas repaired using LaMa[32]. In contrast, Fig. 5(c) clearly illustrates that our method generates more natural and refined character appearances, highlighting the superiority of ImagePD.
Video Comparisons. In the video domain, our approach not only corrects individual frames but also prioritizes inter-frame continuity to mitigate severe visual jitter. As demonstrated in Table 2, VideoPD exhibits superior performance in handling rotational distortions while maintaining optimal stability. In Fig. 6, the red arrows and boxes highlight the unnatural or failed corrections in previous methods when processing video sequences, and these errors typically result in video discontinuity and poor visual perception. In contrast, our framework effectively corrects distortions while preserving the spatiotemporal consistency across the entire video sequence, underscoring the superiority of VideoPD. Additional video results can be found in the supplementary materials.
Method Average Translational Rotational
CoupledTPS[23] 0.9811 0.9785 0.9836
DualPriors[42] 0.9477 0.9290 0.9664
VideoPD(Ours) 0.9886 0.9821 0.9951
Table 2: Quantitative Stability Metrics (Translation/Rotation) for Video Stabilization Models.
Figure 6: Comparison with Existing Portrait Correction Methods in Wide-Angle Videos.
Notably, the impressive performance of our proposed method ImagePD and VideoPD is primarily attributed to the two-stage structural design. This framework combines the transformer’s capacity for long-range dependency modeling with the diffusion model’s multi-step denoising process, achieving global structural consistency and local refinement. By integrating these complementary mechanisms within a unified architecture, our approach achieves remarkable generation quality while maintaining spatial-temporal consistency across various distortion scenarios.
4.4 Ablation studies
In order to demonstrate the effectiveness of our approach and verify the impact of different components, we conducted a series of ablation experiments. In particular, we focus on different diffusion conditions, collaborative architecture design, and video correction trajectories.
Impact of Co-Designed Architecture and Conditional Guidance. Specifically, to independently analyze the impact of each component on correction quality, we designed four ablation scenarios: 1) Transformer Only: The Transformer predicts optical flow without diffusion, testing its global structure modeling. 2)Diffusion with Condition ($I_{s}$): The diffusion model exclusively uses the distorted image $I_{s}$ to predict optical flow, assessing its baseline detail correction. 3) Diffusion with Condition ($I_{s},M$): A mask is applied that marks the face region in the distorted wide-angle image, serving as the conditioning input to guide the model’s generation process. 4) Transformer + Diffusion with Condition ($I_{s},M$): The Transformer provides a feature vector $h$ concatenated with $I_{s}$ and $M$ to guide the diffusion model, evaluating their combined structural and detail refinement capabilities.
Transformer Diffusion $M$ $I_{s}$ LineAcc ShapeAcc
66.617 95.533
66.249 97.465
66.687 97.488
66.898 97.508
Table 3: Ablations on the structure of co-designed architecture and different conditions.
Table  3 presents the experimental results, clearly demonstrating the impact of each module. Specifically, when the Transformer model is integrated with the diffusion model, the scores of LineAcc and ShapeAcc achieve their peak values. These findings confirm our hypothesis that the joint transformer-diffusion architecture can significantly improve the performance of wide-angle lens distortion correction. Furthermore, the progressive incorporation of conditioning information highlights how each component contributes to the robustness and accuracy of portrait stability.
Method Stability LineAcc ShapeACC
ImagePD 0.9325 66.898 97.508
VideoPD 0.9886 66.692 97.480
Table 4: Ablation Experiments Comparing ImagePD and VideoPD on Video Stabilization and Correction.
Impact of Spatiotemporal Smoothing on Video Correction. Our VideoPD model is a fine-tuned extension of the ImagePD model, specifically designed to address the challenges of video correction. Meanwhile, as mentioned before, the key difference between video and photo correction is the spatiotemporal consistency. Therefore, in this section, we directly compare the video results generated by VideoPD and ImagePD to evaluate whether VideoPD alleviates the warping shake caused by the correction process.
In fact, to mitigate the video shaking , some of the original correction of ImagePD will inevitably be compromised. In our experiments, we strive to balance correction and smoothing. This balance ensures that while we reduce the warping shake, we do not sacrifice too much of the correction accuracy. Tab. 4 presents the results. ImagePD achieves excellent correction accuracy, while VideoPD, with spatiotemporal smoothing, increases Stability to 0.9886 (a 6.0% improvement), significantly reducing jitter. The minor decrease in Line/ShapeACC is statistically insignificant ($p>0.05$, t-test), indicating that VideoPD effectively balances correction accuracy and inter-frame smoothness.
For further video results and more detailed analysis, readers can refer to the supplementary materials.
5 Conclusion
In this work, we introduce a novel approach to address facial distortion correction in both images and videos, marking the first application of diffusion models in the domain of portrait distortion correction. Our method represents the first deep learning-based solution for wide-angle video portrait correction, offering a significant advancement over traditional approaches. By leveraging image-level labels to learn the task of unlabeled wide-angle video correction, we drastically reduce the annotation burden and costs. Additionally, our proposed unsupervised framework, VideoPD, enhances spatiotemporal consistency while maintaining a smooth balance between correction and stabilization. We validate our method using a newly created wide-angle video dataset, demonstrating superior performance compared to existing approaches, showcasing its potential for wide-scale, automated, and accurate correction of wide-angle portrait distortion in still images and dynamic video sequences.